OpenStudy (anonymous):

The champion National League baseball team one year one 62 more games than it lost. The team that came second played 154 games winning 16 games less and losing 18 games more than the first team. How many games did each team win and how many did each lose?

OpenStudy (tad1):

w = l + 62 w -16 + l+ 18 = 154 solve the system of equations

OpenStudy (tad1):

-w +L = -62 w + L = 156 2L = 94 L = 47 w = 109 from there you can probably answer the question

OpenStudy (anonymous):

Thanks!!